As a designner you have to obtain results using resistors to follow equation :

R3 <- R1 + R2

What would be an appropriate circuit to do so and why?


1
Expert's answer
2021-09-26T11:55:48-0400

From the understanding that:

The equivalent resistance of the series connection is:Req​=R+R+R+R=4Rohm
Total current in the circuit ,I=V/R=12/4R=3/R Reading of voltmeter A = Voltage across R1​=I×R1​=3/R×R=3V Reading of voltmeter B = Voltage across R2​=I×R2​=3/R×R=3V

 

 I believe that the appropriate circuit to do so is RZ because it maximizes voltage output